OpenStudy (anonymous):

given a polynomial f(x) if (x+3) is a factor what else must be true? f(3)=0 f(-3)=0 f(0)=3 f(0)=-3

OpenStudy (anonymous):

i was thinking a

OpenStudy (kc_kennylau):

Remainder theorem: The remainder of f(x) divided by f(x-a) equals f(a)

OpenStudy (kc_kennylau):

nope :/

OpenStudy (solomonzelman):

f(-3) =0 would be that (-3)+3=0 what f(-3) means plug in -3 for x into f(x)=x+3

OpenStudy (kc_kennylau):

but close

OpenStudy (anonymous):

ohh ok so b

OpenStudy (solomonzelman):

undead, but 3+0 =0?

OpenStudy (anonymous):